One worker takes 20 hours to do a job. A second worker can do the same job in 15 hours. When a third worker is a added it take the three of them 6 hours to do the job. How long would it take the third worker to do the job?
Answer by richwmiller(17219)

6/20+6/15+6/x=1
3/10+2/5+6/x=1
3/10+4/10+6/x=1
7/10+6/x=1
6/x=3/10
x=20